# Learning Racing From an AI Coach: Effects of Multimodal Autonomous Driving Explanations on Driving Performance, Cognitive Load, Expertise, and Trust

### v2

@@ Abstract
In a pre-post experiment (n = 41), we test the impact of an AI Coach’s explanatory communications modeled after the instructions of human driving experts. Participants were divided into four (4) groups to assess two (2) dimensions of the AI coach’s explanations: information type (‘what’ and ‘why’-type explanations) and presentation modality (auditory and visual). We directly compare how AI Coaching sessions employing these techniques impact driving performance, cognitive load, confidence, expertise, and trust in an observation learning context. Through interviews, we delineate the learning process of our participants. Results show that an AI driving coach can be useful for teaching performance driving skills to novices. Comparing between groups, we find the type and modality of information influences performance outcomes. We attribute differences to how information directed attention, mitigated uncertainty, and influenced overload experienced by participants. These, in turn, affected how successfully participants were able to learn. Results suggest efficient, modality-appropriate explanations should be opted for when designing effective HMI communications that can instruct without overwhelming. Further, they support the need to align communications with human learning and cognitive processes. Results are synthesized into eight design implications for future autonomous vehicle HMI and AI coach design.
footnotetext: CC BY: Creative Commons Attribution.
@@ Introduction
Though novel in the realm of autonomous driving, learning from AI, particularly by means of AI explanations, is not a new concept [8]. Learning from AI is a rapidly expanding area of interest, particularly given the proliferation of AI-based large language model tools such as ChatGPT [9, 10]. In domains such as medicine, AI-based systems – for example, image classification systems for radiology – have been shown to help physicians learn new techniques to identify pathologies [11, 12, 13]. Specific cases where human-interpretable AI explanations of decisions are produced have been singled out as especially helpful for learning and improvement [14, 15, 16], particularly in cases where AI explanations are based on the explanations of human experts [17, 18]. AI has the ability to advance educational techniques in many other domains, from second language learning [19] to programming [20].
In the context of autonomous driving, we propose that observing an AI driving coach may be an effective way to transfer knowledge and teach critical driving skills. There have been a number of studies on human-machine interfaces (HMIs) for AVs, which often take the form of Heads-Up-Displays (HUDs) [21]. These studies have majorly focused on driver comprehension [22], trust [23, 24, 25], and explainability [26, 27, 28, 29]. Though vehicle HMIs provide a promising avenue for coaching applications, this instructional application is currently unexplored. The work presented here seeks to contribute to the knowledge gap that exists between HMI design research and learning communities focused on optimal coaching techniques.
@@
To address these gaps of knowledge, this study uses a mixed-methods approach to assess study outcomes. Participants were randomized into one of four AI coaching groups differing by the type and presentation modality of the information they presented. Before and after observation, measures were taken to compare changes in participant driving performance, cognitive load, confidence, expertise, and trust. Each of these factors have been highlighted as important for the development and adoption of HMIs and explainable AI systems more generally [23, 24, 26, 27, 28]. Interviews were conducted to contextualize findings within the larger learning context, including building a more general understanding on concerns with AVs and how these may be mitigated. This broader view of AVs can help illuminate additional roadblocks that must be addressed for successful future adoption of AI driving coaches.
@@
    What is the learning process of a novice learning performance driving, and how can an AI Driving Coach facilitate this learning?
@@
During the AI coaching observation sessions, the vehicle drove autonomously with no user input. The specific model used to control the vehicle was a reinforcement learning agent trained on the specific Thunderhill track. The agent drove in an ideal fashion while considering the physical limitations of the vehicle and track. The RL model was also used as a means to compute a ground truth “ideal” racing line, which could be used to calculate participant performance.
@@
The study was approved for human subjects research by an external IRB and informed consent was collected from all study participants before procedures began. Figure 3 shows the study procedure.
@@ Discussion
1 Simulator calibration
In this phase, participants begin to calibrate the physical inputs on the simulator’s steering wheel and pedals to the movement of the on-screen vehicle.
@@ 4 Vehicle
5 Handling in position
In this phase, participants build their handling skills by iteratively combining their understanding of vehicle positioning with their knowledge of the vehicle’s limits. This is a mediary step between phase 4 with phase 6, and becomes increasingly important as the limits of speed and acceleration are tested. This was the first target for the ‘why’ explanations presented.
@@ 6 Speed via
7 Iteration
As drivers learn, they are able to build on their knowledge, iterate, and enhance specific motor skills, maneuvers, and racing strategies.

### v3

# Effects of Multimodal Explanations for Autonomous Driving on Driving Performance, Cognitive Load, Expertise, Confidence, and Trust

@@ Abstract
Advances in autonomous driving provide an opportunity for AI-assisted driving instruction that directly addresses the critical need for human driving improvement. How should an AI instructor convey information to promote learning? In a pre-post experiment (n = 41), we tested the impact of an AI Coach’s explanatory communications modeled after performance driving expert instructions. Participants were divided into four (4) groups to assess two (2) dimensions of the AI coach’s explanations: information type (‘what’ and ‘why’-type explanations) and presentation modality (auditory and visual). We compare how different explanatory techniques impact driving performance, cognitive load, confidence, expertise, and trust via observational learning. Through interview, we delineate participant learning processes. Results show AI coaching can effectively teach performance driving skills to novices. We find the type and modality of information influences performance outcomes. Differences in how successfully participants learned are attributed to how information directs attention, mitigates uncertainty, and influences overload experienced by participants. Results suggest efficient, modality-appropriate explanations should be opted for when designing effective HMI communications that can instruct without overwhelming. Further, results support the need to align communications with human learning and cognitive processes. We provide eight design implications for future autonomous vehicle HMI and AI coach design.
@@ Introduction
Though novel in the realm of autonomous driving, learning from AI, particularly by means of AI explanations, is not a new concept [8]. Learning from AI is a rapidly expanding area of interest, particularly given the proliferation of AI-based large language model tools such as ChatGPT [9, 10]. In domains such as medicine, explainable AI-based systems – for example, image classification systems for radiology – have been shown to help physicians learn new techniques to identify pathologies [11, 12, 13]. These systems present justifications in the form of ’explanations’ which provide a rationale for a system’s decisions. Specific cases where human-interpretable AI explanations of decisions are produced have been singled out as especially helpful for learning and improvement [14, 15, 16], particularly in cases where AI explanations are based on the explanations of human experts [17, 18]. AI has the ability to advance educational techniques in many other domains, from second language learning [19] to programming [20]. Measuring the impact of interacting with AI systems on learning outcomes can be a challenge, and we can differentiate between explicitly testing learning via knowledge tests and operationalizing learning more functionally via outcomes like task performance [21]. For the present study, we measure learning in both ways. The main learning outcomes assess changes in driving performance before and after exposure to the AI coach; secondarily, we assess knowledge via quiz. In this way, we are able to measure the learning impact of the AI Coach directly from multiple perspectives.
In the context of autonomous driving, we propose that observing an AI driving coach may be an effective way to transfer knowledge and teach critical driving skills. There have been a number of studies on human-machine interfaces (HMIs) for AVs, which often take the form of Heads-Up-Displays (HUDs) aimed at building transparency, accountability, and trust [22, 23]. Though highly relevant, these studies have majorly focused on non-learning areas of the driver experience, and none to our knowledge have tested the impact of HMI elements on driver learning specifically. For example, in the large corpus of work on human-AV trust formation, Morra et al. found that showing more informative displays increased participant willingness to test a real AV [24], Ruijten et al found that displays mimicking human explanations increases trust [25], and Koo et al. found that including why-type information in display improves trust [26]. Other HMI studies have emphasize the complexity of trust formation, proposing holistic and multi-factor approaches to designing trustworthy interfaces [27, 28, 29]. Beyond trust, Schartmüller et al. assessed display modality differences in driver comprehension and multi-tasking, finding that certain interfaces could reduce workload better than others [30]. HMI experiments aimed at building situational awareness with AVs have shown improved awareness with display elements that allow a person to comprehend of why or why-not (contrastive) a vehicle is taking an action [31, 32, 33]. Though these prior studies support vehicle HMIs as a promising avenue for coaching applications, how these may apply in instructional applications is currently unexplored. The work presented here seeks to contribute to the knowledge gap that exists between HMI design research and learning communities focused on optimal coaching techniques.
@@
To address these gaps of knowledge, this study uses a mixed-methods approach to assess study outcomes. Participants were randomized into one of four AI coaching groups differing by the type and presentation modality of the information they presented. Before and after observation, measures were taken to compare changes in participant driving performance, cognitive load, confidence, expertise, and trust. Each of these factors have been highlighted as important for the development and adoption of HMIs and explainable AI systems more generally. Interviews were conducted to contextualize findings within the larger learning context, including building a more general understanding on concerns with AVs and how these may be mitigated. This broader view of AVs can help illuminate additional roadblocks that must be addressed for successful future adoption of AI driving coaches.
@@
    What is the process of a novice learning performance driving, and how can an AI Driving Coach facilitate this learning?
@@
During the AI coaching observation sessions, the vehicle drove autonomously with no user input. The specific model used to control the vehicle was a reinforcement learning agent trained using the DSAC algorithm [40], with observations and rewards tuned for Thunderhill. The agent’s policy was optimized for the physical features and limitations of the vehicle and the geometry of the track [41]. This model was also used to compute the “ideal” racing line for calculating performance.
@@
The study was approved for human subjects research by WCG IRB (external) and informed consent was collected from all study participants before procedures began. All study procedures were performed in accordance with relevant guidelines and regulations. Figure 3 shows the study procedure.
@@ Discussion
1 Simulator calibration In this phase, participants begin to calibrate the physical inputs on the simulator’s steering wheel and pedals to the movement of the on-screen vehicle.
@@ 4 Vehicle
5 Handling in position In this phase, participants build their handling skills by iteratively combining their understanding of vehicle positioning with their knowledge of the vehicle’s limits. This is a mediary step between phase 4 with phase 6, and becomes increasingly important as the limits of speed and acceleration are tested. This was the first target for the ‘why’ explanations presented.
@@ 6 Speed via
7 Iteration As drivers learn, they are able to build on their knowledge, iterate, and enhance specific motor skills, maneuvers, and racing strategies.